Assume that the banking system is loaned up and that any open-market purchase by the Fed directly increases reserves in the banks. If the required reserve ratio is 0, by how much could the money supply expand if the Fed purchased $2 billion worth of bonds?


1
Expert's answer
2021-12-20T17:57:39-0500

Solution:

First, derive the money multiplier:

Money multiplier = "\frac{1}{r} = \frac{1}{0} = 0"


Change in money supply = Change in money reserves "\times" the multiplier


Change in money supply = 2 "\times" 0 = 0

The money supply will not expand or change if the required reserve ratio is equal to zero.